OpenStudy (anonymous):

The price of a video game is $32.99 before tax. Andre bought the video game for 20% off. He then paid 8% sales tax on the discounted price. Part A: How much did Andre pay in sales tax? Show all work and steps in your solution. Part B: What is the total amount that Andre paid for the video game? Show all work and steps in your solution. I got 32.99 * 20/100=6.60. 32.99-6.60= 26.39. 26.39* 8/100=2.11. 26.39+2.11=28.50?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@kidrah69

OpenStudy (anonymous):

so my answer would be: A: $2.11 B: $28.50 Work: 32.99 * 20/100=6.60. 32.99-6.60= 26.39. 26.39* 8/100=2.11. 26.39+2.11=28.50

OpenStudy (anonymous):

right?

OpenStudy (kidrah69):

wheres the work for part A?

OpenStudy (anonymous):

its in the work.

OpenStudy (anonymous):

the tax would be $2.11

OpenStudy (anonymous):

A: $2.11 B: $28.50 Work: 32.99 * 20/100=6.60. 32.99-6.60= 26.39. 26.39* 8/100=2.11. 26.39+2.11=28.50

OpenStudy (kidrah69):

look right

OpenStudy (anonymous):

ok thanks

OpenStudy (kidrah69):

:)